# Unified description of viscous, viscoelastic, or elastic thin active films on substrates

### v1

@@ Appendix A Elastic stress and displacement field
$\delta_{ij}$ denotes the Kronecker delta, $\mu$ is the shear modulus, and the material parameter $\lambda$ diverges for incompressible materials. We here confine ourselves to orders linear in strain, that is, linear in gradients of the displacement field $\nabla\mathbf{u}$. Thus, the components of the strain tensor reduce to
@@
In our considerations, we assume the materials to be incompressible. The diverging material parameter $\lambda$ leads to a vanishing trace of $\mathbf{U}$ to linear order, that is, $\nabla\cdot\mathbf{u}=0$. Ensuring that this condition is satisfied allows us to drop the contribution of $\lambda$ in the following.
@@
$d\epsilon=\Psi_{ij}dU_{ij}\,$ (48)
allows us to write the components of the associated thermodynamic conjugate $\boldsymbol{\Psi}$ as
@@
The elastic shear forces enter Eq. (18) through the divergence of the elastic stress tensor $\boldsymbol{\sigma}^{\mathrm{el}}$, which reads [59]
$\boldsymbol{\sigma}^{\mathrm{el}}_{ij}=\Psi_{ij}-\Psi_{ki}U_{jk}-\Psi_{kj}U_{ik}\,.$ (50)
We here confine ourselves to the linear form
$\boldsymbol{\sigma}^{\mathrm{el}}_{ij}=\Psi_{ij}=2\mu U_{ij}\,.$ (51)
@@
$\sigma_{ij}^{\mathrm{el}}=\mu(\partial_{x_{i}}u_{j}+\partial_{x_{j}}u_{i})\,.$ (52)
To close the system of equations, we add the dynamic equation for the strain tensor $\mathbf{U}$ [59],
$\partial_{t}\mathbf{U}+\mathbf{v}\cdot\nabla\mathbf{U}+(\nabla\mathbf{v})\cdot\mathbf{U}+\mathbf{U}\cdot(\nabla\mathbf{v})^{\top}=\boldsymbol{\Sigma}-\frac{1}{\tau_{\mathrm{d}}}\mathbf{U}\,.$ (53)
It includes as the second term on the left-hand side the translational convective transport with the flow field $\mathbf{v}$. Rotational convective transport is described by the subsequent two terms, where marks the transpose. On the right-hand side, strains are generated by flows according to the deformation rate tensor $\boldsymbol{\Sigma}=[(\nabla\mathbf{v})^{\top}+(\nabla\mathbf{v})]/2$. Up to this term, the form of the equation can be rigorously derived from conservation laws and symmetry arguments [59]. We again assume isotropic elasticity as well as incompressibility, that is, $U_{ii}=0$. An additional relaxation term enters the right-hand side in the case of viscoelastic systems that are not strictly elastic solid-like [59]. In our framework, it describes the loss of memory of previous undeformed states over time. Thus, when elastic stresses are released, the material elements are not fully displaced back to their initial positions, but to some intermediate locations. The material has flown a certain distance. $\tau_{\mathrm{d}}$ denotes the corresponding relaxation time.
We now insert into Eq. (53) the expression for the strain tensor, Eq. (47). Accordingly, the field $\mathbf{u}$ is interpreted in an Eulerian framework as well. Thus, it describes the memory of the reversible elastic displacements that are currently stored in the material and have taken the material elements to their current positions $\mathbf{x}$ at time $t$. In other words, if all stresses were switched off at time $t$, a material element currently located at position $\mathbf{x}$ tends to move back to $\mathbf{x}-\mathbf{u}(\mathbf{x},t)$. From Eq. (53), we obtain
@@
$+(\partial_{x_{j}}v_{k})(\partial_{x_{i}}u_{k})+(\partial_{x_{j}}v_{k})(\partial_{x_{k}}u_{i})=\partial_{x_{i}}v_{j}+\partial_{x_{j}}v_{i}-\frac{1}{\tau_{\mathrm{d}}}(\partial_{x_{i}}u_{j}+\partial_{x_{j}}u_{i})\,.$
We aim for a consistent theory in the incompressible limit, that is, $\nabla\cdot\mathbf{v}=\partial_{x_{i}}v_{i}=0$ and $\nabla\cdot\mathbf{u}=\partial_{x_{i}}u_{i}=0$ must be fulfilled at all times. Taking the trace of Eq. (54) yields an evolution equation for $\nabla\cdot\mathbf{u}$,
$\partial_{t}\partial_{x_{i}}u_{i}+v_{k}\partial_{x_{k}}\partial_{x_{i}}u_{i}+(\partial_{x_{i}}v_{k})(\partial_{x_{k}}u_{i})+(\partial_{x_{i}}v_{k})(\partial_{x_{i}}u_{k})=-\frac{1}{\tau_{\mathrm{d}}}\partial_{x_{i}}u_{i}\,,$ (55)
@@
$\partial_{t}\partial_{x_{i}}u_{j}+\partial_{t}\partial_{x_{j}}u_{i}+v_{k}\partial_{x_{k}}\partial_{x_{i}}u_{j}+v_{k}\partial_{x_{k}}\partial_{x_{j}}u_{i}=\partial_{x_{i}}v_{j}+\partial_{x_{j}}v_{i}-\frac{1}{\tau_{\mathrm{d}}}(\partial_{x_{i}}u_{j}+\partial_{x_{j}}u_{i})\,.$ (56)
@@
The left-hand side is consistent with Eq. (20) for vanishing right-hand side of the equation. Equation (20) only contains couplings between the velocity itself (not its gradient) to gradients of the displacement field. On the right-hand side of Eq. (59), we find the nonlinear couplings between velocity gradients and gradients of the displacement field, as well as the auxiliary field $q$. Overall, we accept deviations from nonlinear terms that result from our confinement to linear elasticity under couplings to the velocity field. However, the description shall maintain our condition of incompressibility $\nabla\cdot\mathbf{u}=0$ within the linear regime. To this end, we take the trace of Eq. (59) and enforce its right-hand side to vanish, which leads to an equation for the auxiliary field $q$,
@@
It serves to determine the field $q$ and together with Eq. (20) describes the evolution of the displacement field $\mathbf{u}$ in a consistent way when confining ourselves to linear order in $\nabla\mathbf{u}$ and incompressible systems.
Physically, these considerations imply that the nonlinear term associated with translational convection of the displacement field is maintained. This is one step beyond ordinary linear elasticity. In particular, this contribution becomes significant when flows arise for viscoelastic fluids and therefore is physically significant.
@@ Appendix C Numerical methods
We use a pseudo-spectral method [61] to solve Eqs. (23)–(25) in a two-dimensional domain under periodic boundary conditions. This scheme exploits the fast Fourier transformation to calculate spatial derivatives, which significantly accelerates solving Eq. (24). Time integration is performed via a fourth-order Runge–Kutta method using a time step of $\Delta t=10^{-3}$ in our rescaled units. To ensure that the velocity field $\mathbf{v}$ stays divergence-free, we use a projection method [62]. The pressure $p$ is determined in a way to satisfy the incompressibility conditions. If not stated otherwise, our numerical calculations start from a weakly perturbed isotropic state. Random initial values taken from a uniform distribution over the interval $[-0.01,0.01]$ are assigned to each of the components of the polar orientational order parameter field at every grid point. Velocity and displacement fields are set to zero initially. When investigating hysteretic behavior, we successively change the activity as quantified by the parameter $\nu_{\mathrm{p}}$, see Eq. (19), and use the results of the previous calculation as a starting point for the next step. To obtain the data shown in Fig. 2 and Fig 3(a), we use system sizes of $5\times 5$ in rescaled units, spanned by a numerical grid of $32\times 32$ points. We analyze the dynamics in a window of $1000$ time units to measure each of the displayed data points Fig 3(a).

### v2

@@ Appendix A Elastic stress and memory field of elastic displacements
$\delta_{ij}$ denotes the Kronecker delta, $\mu$ is the shear modulus, and the material parameter $\lambda$ diverges for incompressible materials. We here confine ourselves to orders linear in strain, that is, linear in gradients $\nabla\mathbf{u}$. Thus, the components of the strain tensor reduce to
@@
In our considerations, we assume the materials to be incompressible. The diverging material parameter $\lambda$ leads to a vanishing trace of $\mathbf{U}$ to linear order, that is, $\nabla\cdot\mathbf{u}=0$. Ensuring that this condition remains satisfied, see below, allows us to drop the contribution of $\lambda$ in the following.
@@
$d\epsilon=\Psi_{ij}dU_{ij}\,,$ (47)
allows us to write the components of the associated thermodynamic conjugate $\bm{\Psi}$ as
@@
The elastic shear forces enter the flow equation for the velocity field $\mathbf{v}$, Eq. (18), through the divergence of the elastic stress tensor $\bm{\sigma}^{\mathrm{el}}$, see also the paragraph below Eq. (18). The components of $\bm{\sigma}^{\mathrm{el}}$ read [62]
${\sigma}^{\mathrm{el}}_{ij}=\Psi_{ij}-\Psi_{ki}U_{jk}-\Psi_{kj}U_{ik}\,.$ (49)
We here confine ourselves to terms linear in elastic strain, so that
${\sigma}^{\mathrm{el}}_{ij}=\Psi_{ij}=2\mu U_{ij}\,.$ (50)
@@
$\sigma_{ij}^{\mathrm{el}}=\mu(\partial_{x_{i}}u_{j}+\partial_{x_{j}}u_{i})\,$ (51)
in terms of the memory field of elastic displacements $\mathbf{u}$. Due to incompressibility, $\nabla\cdot\bm{\sigma}^{\mathrm{el}}=\mu\nabla^{2}\mathbf{u}$, which enters the flow equation for $\mathbf{v}$, Eq. (18).
Second, the dynamic part of the generalized hydrodynamics theory associated with (visco)elasticity [62, 67] is the dynamic equation for the strain tensor. In its vectorial notation, that is, for the elastic memory field $\mathbf{u}$, it corresponds to our Eq. (20). To begin, this dynamic equation in Ref. 62 linearized in strain $\mathbf{U}$ reads
$\partial_{t}\mathbf{U}+\mathbf{v}\cdot\nabla\mathbf{U}+(\nabla\mathbf{v})\cdot\mathbf{U}+\mathbf{U}\cdot(\nabla\mathbf{v})^{\top}=\bm{\Sigma}-\tau_{\mathrm{d}}^{-1}\mathbf{U}\,.$ (52)
It includes as the second term on the left-hand side the translational convective transport with the flow field $\mathbf{v}$. Lower convective transport is described by the subsequent two terms [67], where marks the transpose. On the right-hand side, strains are generated by flows, that is, motion of material, according to the deformation rate tensor $\bm{\Sigma}=[(\nabla\mathbf{v})^{\top}+(\nabla\mathbf{v})]/2$. The form of the equation can be rigorously derived from conservation laws and symmetry arguments [62]. Without the relaxation term ${}-\tau_{\mathrm{d}}^{-1}\mathbf{U}$, Eq. (52) describes genuinely elastic systems. We again assume isotropic elasticity as well as incompressibility, that is, $U_{ii}=0$. An additional relaxation term, here denoted as ${}-\tau_{\mathrm{d}}^{-1}\mathbf{U}$, enters the right-hand side in the case of viscoelastic systems that are not strictly elastic solid-like [62]. In our framework, it describes the loss of memory about previous undeformed states over time. Thus, when elastic stresses are released, the material elements are not fully displaced back to their initial positions, but to some intermediate locations. In other words, the material has flown a certain distance. $\tau_{\mathrm{d}}$ denotes the corresponding relaxation time.
We now insert into Eq. (52) the expression for the linearized strain tensor, Eq. (46). Accordingly, the field $\mathbf{u}$ is interpreted in an Eulerian framework as well. Thus, it describes the memory of the reversible elastic displacements that are presently stored in the material for the material elements currently located at positions $\mathbf{x}$ at time $t$. In other words, if all imposed stresses were switched off at time $t$, the material element currently located at position $\mathbf{x}$ at time $t$ tends to move back to $\mathbf{x}-\mathbf{u}(\mathbf{x},t)$. From Eq. (52), we obtain
@@
$+(\partial_{x_{j}}v_{k})(\partial_{x_{i}}u_{k})+(\partial_{x_{j}}v_{k})(\partial_{x_{k}}u_{i})=\partial_{x_{i}}v_{j}+\partial_{x_{j}}v_{i}-{\tau_{\mathrm{d}}}^{-1}(\partial_{x_{i}}u_{j}+\partial_{x_{j}}u_{i})\,.$
We aim for a consistent theory in the incompressible limit, that is, $\nabla\cdot\mathbf{v}=\partial_{x_{i}}v_{i}=0$ and $\nabla\cdot\mathbf{u}=\partial_{x_{i}}u_{i}=0$ must be satisfied at all times. Taking the trace of Eq. (53) yields an evolution equation for $\nabla\cdot\mathbf{u}$,
$\partial_{t}\partial_{x_{i}}u_{i}+v_{k}\partial_{x_{k}}\partial_{x_{i}}u_{i}+(\partial_{x_{i}}v_{k})(\partial_{x_{k}}u_{i})+(\partial_{x_{i}}v_{k})(\partial_{x_{i}}u_{k})=-{\tau_{\mathrm{d}}}^{-1}\partial_{x_{i}}u_{i}\,,$ (54)
@@
$\partial_{t}\partial_{x_{i}}u_{j}+\partial_{t}\partial_{x_{j}}u_{i}+v_{k}\partial_{x_{k}}\partial_{x_{i}}u_{j}+v_{k}\partial_{x_{k}}\partial_{x_{j}}u_{i}=\partial_{x_{i}}v_{j}+\partial_{x_{j}}v_{i}-{\tau_{\mathrm{d}}}^{-1}(\partial_{x_{i}}u_{j}+\partial_{x_{j}}u_{i})\,.$ (55)
@@
For vanishing right-hand side, the left-hand side is consistent with Eq. (20). Equation (20) only contains couplings between the velocity itself (not its gradient) and gradients of the memory field of elastic displacements $\mathbf{u}$. On the right-hand side of Eq. (58), we find nonlinear couplings between velocity gradients and gradients of the elastic memory field $\mathbf{u}$, as well as the auxiliary field $q$. Since we confine ourselves to linear elasticity, meaning linearity in gradients of $\mathbf{u}$, the description shall maintain our condition of incompressibility $\nabla\cdot\mathbf{u}=0$ within the linear regime. To this end, we take the trace of Eq. (58) and enforce its right-hand side to vanish, which leads to an equation for the auxiliary field $q$,
@@
It serves to determine the field $q$ and together with Eq. (20) describes the evolution of the elastic memory field $\mathbf{u}$ in a consistent way when confining ourselves to linear order in $\nabla\mathbf{u}$ and incompressible systems.
Physically, these considerations imply that the nonlinear term associated with translational convection of the elastic memory field $\mathbf{u}$ is maintained. This is one step beyond ordinary linear elasticity. In particular, this contribution becomes important when flows arise for viscoelastic fluids and therefore is physically significant.
@@ Appendix C Numerical methods
We use a pseudo-spectral method [88] to solve Eqs. (22)–(24) in a two-dimensional domain under periodic boundary conditions. This scheme exploits the fast Fourier transformation to calculate spatial derivatives, which significantly accelerates solving Eq. (23). Time integration is performed via a fourth-order Runge–Kutta method using a time step of $\Delta t=10^{-3}$ in our rescaled units. To ensure that the velocity field $\mathbf{v}$ stays divergence-free, we use a projection method [89]. The pressure $p$ is determined in a way to satisfy the incompressibility conditions. If not stated otherwise, our numerical calculations start from a weakly perturbed isotropic state. Random initial values taken from a uniform distribution over the interval $[-0.01,0.01]$ are assigned to each of the components of the polar orientational order parameter field at every grid point. Velocity and elastic memory fields are set to zero initially. When investigating hysteretic behavior, we successively change the activity as quantified by the parameter $\nu_{\mathrm{p}}$, see Eq. (19), and use the results of the previous calculation as a starting point for the next step. To obtain the data shown in Fig. 2 and Fig 3(a), we use system sizes of $5\times 5$ in rescaled units, spanned by a numerical grid of $32\times 32$ points. We analyze the dynamics in a window of $1000$ time units to measure each of the displayed data points in Fig 3(a).